# FFCBA: Feature-based Full-target Clean-label Backdoor Attacks

### v1

@@ Abstract.
Backdoor attacks pose a significant threat to deep neural networks, as backdoored models would misclassify poisoned samples with specific triggers into target classes while maintaining normal performance on clean samples. Among these, multi-target backdoor attacks can simultaneously target multiple classes. However, existing multi-target backdoor attacks all follow the dirty-label paradigm, where poisoned samples are mislabeled, and most of them require an extremely high poisoning rate. This makes them easily detectable by manual inspection. In contrast, clean-label attacks are more stealthy, as they avoid modifying the labels of poisoned samples. However, they generally struggle to achieve stable and satisfactory attack performance and often fail to scale effectively to multi-target attacks. To address this issue, we propose the Feature-based Full-target Clean-label Backdoor Attacks (FFCBA) which consists of two paradigms: Feature-Spanning Backdoor Attacks (FSBA) and Feature-Migrating Backdoor Attacks (FMBA). FSBA leverages class-conditional autoencoders to generate noise triggers that align perturbed in-class samples with the original category’s features, ensuring the effectiveness, intra-class consistency, inter-class specificity and natural-feature correlation of triggers. While FSBA supports swift and efficient attacks, its cross-model attack capability is relatively weak. FMBA employs a two-stage class-conditional autoencoder training process that alternates between using out-of-class samples and in-class samples. This allows FMBA to generate triggers with strong target-class features, making it highly effective for cross-model attacks. We conduct experiments on multiple datasets and models, the results show that FFCBA achieves outstanding attack performance and maintains desirable robustness against the state-of-the-art backdoor defenses.
@@
ccs: Security and privacyccs: Computing methodologies Computer vision
@@ 1. Introduction
  • We propose FFCBA, consisting of FSBA and FMBA. Both achieve low-poisoning-rate clean-label full-target backdoor attacks. FSBA is more efficient and faster, while FMBA has excellent cross-model attack capability.
@@
  • Our experiments confirm that FFCBA exhibits strong attack capabilities, with minimal impact on benign accuracy and desirable robustness against advanced backdoor defenses.
@@ 2. Related Work
2.1. Backdoor Attacks
Dirty-label Backdoor Attacks. BadNets (Gu et al., 2019) first unveiled the backdoor attack chapter, with subsequent studies (Chen et al., 2017; Liu et al., 2020; Nguyen and Tran, 2021; Feng et al., 2022; Wang et al., 2022; Gao et al., 2024b) enhancing stealth and potency. However, these efforts concentrated on single-target attacks, resulting in limited attack capabilities. Works like One-to-N (Xue et al., 2020), Marksman (Doan et al., 2022), and Universal Backdoor (Schneider et al., 2024) break this limitation achieving multi-target attacks. However, they all operate under dirty label conditions and, except for Universal Backdoor (Schneider et al., 2024), require extremely high poisoning rates. Thus, they struggle to evade human detection, often leading to failure. Addressing the challenge of conducting low-poisoning-rate clean-label multi-target attacks is an urgent issue that requires resolution.
@@ 2.1. Backdoor Attacks
2.2. Backdoor Defenses
A detailed introduction to backdoor defenses is in Appendix A.1. The complete appendix is provided in the supplementary material.
@@ 3. Threat Model
[…]
@@ 4.2. FSBA Paradigm
(1) Output Layer Loss. We aim for the mixed samples $x_{m,k}$ to be classified as their original category $y_{k}$ by the proxy classification model $f_{c}$. This enables the trigger $T_{k}$ to suppress the mid-high-frequency perturbations in $x_{m,k}$, allowing $x_{m,k}$ to exhibit the features of the original category $y_{k}$. Therefore, the trigger-feature strength $>$ perturbation strength $>$ natural-feature strength, ensuring the trigger effectiveness. Additionally, triggers for each class can exhibit robust features intrinsic to that class, guaranteeing the inter-class specificity and the natural-feature correlation. To accomplish the aim, we formulate the output layer loss function as:
@@
(2) Latent Space Loss. Relying solely on ensuring that the mixed samples $x_{m,k}$ fall within the classification boundary of their original category $y_{k}$ poses challenges in meeting the intra-class consistency of the trigger. Specifically, in datasets with limited categories, the classification boundaries are often quite loose. This results in dispersed trigger feature distributions targeting the same class, leading to insufficient consistency. To address the issue, we calculate the centroids of the feature vector clusters for each category in the latent space of $f_{c}$, denoted as $Mean$. Here, the latent space refers to the representation of inputs in the penultimate layer of the model, denoted as $\mathcal{Z}$. Consequently, $f_{c}$ can be divided into two parts: the feature extraction component $z_{c}:X\rightarrow\mathcal{Z}$ and the linear classification component $l_{c}:\mathcal{Z}\rightarrow C$, where $f_{c}=z_{c}\circ l_{c}$. This means that the classification result of $f_{c}$ is achieved by first applying $z_{c}$, followed by $l_{c}$. By constraining the distribution of the mixed sample $x_{m,k}$ in the latent space to cluster around the centroid $Mean(y_{k})$ of the original category $y_{k}$, we can significantly reduce the dispersion of the noise trigger features. Thus, we can derive the latent space loss as 6, where $\mathcal{L}_{1}$ denotes the L1 loss.
@@ 4.3. FMBA Paradigm
where $x_{k,i}$, and $v_{k}$ represent the training samples, and the one-hot label vectors of category $y_{k}$, respectively. $K$ and $n_{k}$ have the same meanings as in Eq. (5). The output is a vector with same dimension as $v_{k}$, representing the probabilities of being classified into each category. Following EBBA (Gao et al., 2024a), $K(x,x_{k,i})=e^{-2\gamma||x-x_{k,i}||^{2}}$, $\gamma>0$. The ratio of the probabilities that sample $x_{a}$ being classified into category $y_{a}$ and $y_{b}$ by $f$ tends toward a fixed multiple $\lambda$ as:
(10) $\begin{split}\frac{\sum_{i=0}^{n_{a}}e^{-2\gamma||x_{a,i}||^{2}}e^{4\gamma x_{a}\cdot x_{a,i}}}{\sum_{i=0}^{n_{b}}e^{-2\gamma||x_{b,i}||^{2}}e^{4\gamma x_{a}\cdot x_{b,i}}}=\lambda.\end{split}$
@@
(11) $\begin{split}\frac{\sum_{i=0}^{n_{a}}e^{-2\gamma||x_{a,i}||^{2}}e^{4\gamma x_{a}\cdot x_{a,i}}e^{4\gamma x_{b}\cdot x_{a,i}}}{\sum_{i=0}^{n_{b}}e^{-2\gamma||x_{b,i}||^{2}}e^{4\gamma x_{a}\cdot x_{b,i}}e^{4\gamma x_{b}\cdot x_{b,i}}}\approx 1.\end{split}$
@@
5. Evaluation
5.1. Experimental Setup
Hyperparameters. The specific parameter settings are detailed in Appendix A.5.
@@ 5.2. Effectiveness of FFCBA
Attack Effectiveness. Table 2 contrasts FFCBA’s attack performance with baselines. FSBA and FMBA, using class-conditional autoencoders trained on the Resnet18 proxy model, target the Resnet18 victim model and execute cross-model attacks on VGG19. We outline the average ASR across all labels, Benign classification Accuracy (BA), and the Decrease Value in BA (DV) compared to the clean model. FFCBA attains high ASR across all labels and datasets while minimally impacting BA, showing advantages over various baselines. Against One-to-N, FFCBA gains a significant ASR boost. Compared to Marksman and Universal Backdoor, FFCBA matches their attack effectiveness under stricter clean-label constraints. Moreover, FFCBA maintains a low 0.4% poisoning rate across all datasets, much lower than Marksman (10%) and One-to-N (20%), and comparable to Universal Backdoor (0.62%). Figure 4 specifically illustrates how the ASR of FFCBA changes with varying poisoning rates. In addition, FFCBA’s poisoned samples exhibit better visual quality than baselines, as illustrated in Appendix A.2. Table 3 presents FFCBA’s ASR for each CIFAR10 label with Resnet18 victim model. FFCBA demonstrates robust attack performance in all categories. Given the large number of categories in Animals90 and ImageNet100, we omit ASR for each label individually.
@@
Cross-model Attack Capability. We further evaluate the cross-model attack capabilities of FFCBA under various conditions. Table 4 presents FFCBA’s cross-model attack performance across different datasets. Specifically, we train class-conditional autoencoders on various datasets using the Resnet18 proxy model to attack other models with significant architectural and scale differences. The results indicate that FMBA consistently exhibits excellent cross-model attack capabilities on any dataset, while FSBA’s performance declines as architectural disparities increase. Table 5 presents FFCBA’s cross-model attack performance across different proxy models. Specifically, we train class-conditional autoencoders on ImageNet100 using Densenet121 and ViT_B_16 proxy models, which have significant structural differences, to launch attacks on various models. The results show that FMBA maintains superior cross-model attack capabilities regardless of the proxy model used, whereas FSBA performs weaker. Despite its limited cross-model capabilities, FSBA can execute swift and potent attacks when the victim model type is known. Table 6 illustrates FSBA’s effectiveness with an accurate proxy model, achieving superior attack outcomes across various model architectures and datasets. Therefore, with knowledge of the victim model type, FSBA can be deployed for quick and efficient attacks. Conversely, when the victim model type is unknown or it is challenging to train an accurate proxy model, FMBA can be effectively employed for cross-model attacks. Their application scenarios are complementary.
5.3. Robustness against Backdoor Defenses
In this section, we evaluate the robustness of FFCBA against popular backdoor defense mechanisms, including Fine-Pruning (Liu et al., 2018a), Neural Cleanse (Wang et al., 2019), STRIP (Gao et al., 2019), CBD (Zhang et al., 2023), EBBA (Gao et al., 2024a), ABL (Li et al., 2021b), and IBD-PSC (Hou et al., 2024). These defense mechanisms have proven to be effective against previous backdoor attacks.
@@ 5.3. Robustness against Backdoor Defenses
Resistance to Fine-Pruning. Fine-Pruning neutralizes backdoors by pruning dormant neurons while ensuring benign accuracy. Figure 5(a) and Figure 5(b) show the resistance of FFCBA to Fine-Pruning. It can be observed that under different pruning rounds, the degree of ASR decline is always lower than BA, hence Fine-Pruning cannot defend against these two attack paradigms.
Resistance to Neural Cleanse. Neural Cleanse detects backdoors by constructing reverse-engineered triggers and measuring whether anomaly metric exceed a threshold of 2. Figure 5(c) illustrates the resistance of FSBA and FMBA to Neural Cleanse. Both exhibit anomaly metrics below the threshold, indicating that Neural Cleanse cannot defend against these two attack paradigms.
Resistance to STRIP. STRIP perturbs clean images to generate high-entropy outputs in benign models. Low entropy indicates the presence of backdoors. Figure 5(e) and Figure 5(f) show that the entropy distributions of clean and poisoned samples of Animals90 for FSBA and FMBA are similar, suggesting that STRIP cannot defend against these two attack paradigms. The results on the other two datasets are provided in Appendix A.6.
Resistance to CBD. CBD learns causal relationships and employs a per-sample weighting scheme on contaminated datasets to obtain a clean model. Figure 5(d) presents the results of FFCBA’s resistance to CBD. It is evident that both FSBA and FMBA exhibit high ASR, demonstrating their robustness against CBD.
Resistance to EBBA. EBBA calculates the energy of each label in the model’s output. If any label exhibits an abnormally high energy value, it suggests the presence of a backdoor. Figure 5(g) and Figure 5(h) present the results of FSBA and FMBA against EBBA on ImageNet100, showing that the energy distribution across labels is quite uniform. The results on the other two datasets are provided in Appendix A.7. Since every label is the target label, EBBA is destined to fail in detecting FFCBA.
Resistance to ABL. ABL isolates backdoor examples at the early training stage and later breaks their correlation with the target class, so that training train clean models based on poisoned data. Figure 5(i) shows the results of FSBA and FMBA against ABL. Both maintain high ASR, indicating they can bypass ABL.
Resistance to IBD-PSC. IBD-PSC detects backdoor attacks by amplifying batch normalization parameters in models, enhancing the prediction confidence consistency of poisoned samples, thus identifying malicious inputs effectively. Figure 5(j) shows the results of FSBA and FMBA against IBD-PSC. Their low AUROC and F1 scores indicate that IBD-PSC cannot defend against FFCBA.

### v2

@@ Abstract.
Backdoor attacks pose a significant threat to deep neural networks, as backdoored models would misclassify poisoned samples with specific triggers into target classes while maintaining normal performance on clean samples. Among these, multi-target backdoor attacks can simultaneously target multiple classes. However, existing multi-target backdoor attacks all follow the dirty-label paradigm, where poisoned samples are mislabeled, and most of them require an extremely high poisoning rate. This makes them easily detectable by manual inspection. In contrast, clean-label attacks are more stealthy, as they avoid modifying the labels of poisoned samples. However, they generally struggle to achieve stable and satisfactory attack performance and often fail to scale effectively to multi-target attacks. To address this issue, we propose the Feature-based Full-target Clean-label Backdoor Attacks (FFCBA) which consists of two paradigms: Feature-Spanning Backdoor Attacks (FSBA) and Feature-Migrating Backdoor Attacks (FMBA). FSBA leverages class-conditional autoencoders to generate noise triggers that align perturbed in-class samples with the original category’s features, ensuring the effectiveness, intra-class consistency, inter-class specificity and natural-feature correlation of triggers. While FSBA supports swift and efficient attacks, its cross-model attack capability is relatively weak. FMBA employs a two-stage class-conditional autoencoder training process that alternates between using out-of-class samples and in-class samples. This allows FMBA to generate triggers with strong target-class features, making it highly effective for cross-model attacks. We conduct experiments on multiple datasets and models, the results show that FFCBA achieves outstanding attack performance and maintains desirable robustness against the state-of-the-art backdoor defenses. Our source code is available at (https://github.com/YangxvYin/FFCBA_code).
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755227isbn: 979-8-4007-2035-2/2025/10ccs: Security and privacyccs: Computing methodologies Computer vision
@@ 1. Introduction
[…]
Schematic of low-poisoning-rate clean-label multi-target backdoor attacks during training and inference phases.
@@
  • FFCBA consists of FSBA and FMBA enabling low-poisoning clean-label full-target attacks. FSBA is more efficient, while FMBA has excellent cross-model attack capability.
@@
  • Experiments show FFCBA delivers strong attacks with negligible benign accuracy impact and robust defense resistance.
@@ 2. Related Work
Dirty-label Backdoor Attacks. BadNets (Gu et al., 2019) first unveiled the backdoor attack chapter, with subsequent studies (Chen et al., 2017; Liu et al., 2020a; Nguyen and Tran, 2021; Feng et al., 2022; Wang et al., 2022; Gao et al., 2024b) enhancing stealth and potency. However, these efforts concentrated on single-target attacks, resulting in limited attack capabilities. Works like One-to-N (Xue et al., 2020), Marksman (Doan et al., 2022), and Universal Backdoor (Schneider et al., 2024) break this limitation achieving multi-target attacks. However, they all operate under dirty label conditions and, except for Universal Backdoor (Schneider et al., 2024), require extremely high poisoning rates. Thus, they struggle to evade human detection, often leading to failure. Addressing the challenge of conducting low-poisoning-rate clean-label multi-target attacks is an urgent issue that requires resolution.
@@
Backdoor Defenses. Current backdoor defense paradigms (Selvaraju et al., 2017; Zhang et al., 2023; Liu et al., 2020b; Chen et al., 2018, 2019; Guan et al., 2024) can be mainly categorized into input-based defenses, model-based defenses, output-based defenses, and inhibition-based defenses. Input-based defenses detect or disrupt triggers in samples to neutralize backdoors. For example, frequency-based detector (Zeng et al., 2021) identifies triggers through high-frequency artifacts. Model-based defenses detect or eliminate backdoor by analyzing or adjusting model parameters. For example, Neural Cleanse (Wang et al., 2019) obtains possible triggers by reverse engineering each class and comparing them to find potential attack targets. Fine-Pruning (Liu et al., 2018a) prunes the model to destroy backdoor structures. Output-based defenses detect the backdoor by checking model outputs. STRIP (Gao et al., 2019) perturbs potentially poisoned images with random clean images and monitors output entropy for backdoor signs. EBBA (Gao et al., 2024a) calculates output energy per class, seeking abnormal high-energy values indicative of backdoors. Inhibition-based defenses aim to train clean models based on poisoned datasets. For example, CBD (Zhang et al., 2023) learns backdoor-free models directly from contaminated datasets from a causal point of view. However, these defense mechanisms struggle to detect FFCBA.
@@ 3. Threat Model
(2) $\small\underset{\theta}{\min}\sum_{i=1}^{N_{b}}\mathcal{L}(f^{\prime}(x_{i};\theta),y_{i})+\sum_{t=1}^{K}\sum_{j=1}^{\lfloor N_{p}/K\rfloor}\mathcal{L}(f^{\prime}(B_{t}(x_{t}^{j});\theta),y_{t}),$
@@ 4.2. FSBA Paradigm
[…]
Schematic of feature changes in poisoned samples during the FFCBA attack process.
@@
(1) Output Layer Loss. We aim for the mixed samples $x_{m,k}$ to be classified as their original category $y_{k}$ by the proxy classification model $f_{c}$. This enables the trigger $T_{k}$ to suppress the mid-high-frequency perturbations in $x_{m,k}$, allowing $x_{m,k}$ to exhibit the features of the original category $y_{k}$. Therefore, the trigger-feature strength >> perturbation strength >> natural-feature strength, ensuring the trigger effectiveness. Additionally, triggers for each class can exhibit robust features intrinsic to that class, guaranteeing the inter-class specificity and the natural-feature correlation. To accomplish the aim, we formulate the output layer loss function as:
@@
(2) Latent Space Loss. Relying solely on ensuring that the mixed samples $x_{m,k}$ fall within the classification boundary of their original category $y_{k}$ poses challenges in meeting the intra-class consistency of the trigger. Specifically, in datasets with limited categories, the classification boundaries are often quite loose. This results in dispersed trigger feature distributions targeting the same class, leading to insufficient consistency. To address the issue, we calculate the centroids of the feature vector clusters for each category in the latent space of $f_{c}$, denoted as $Mean$. Here, the latent space refers to the representation of inputs in the penultimate layer of the model, denoted as $\mathcal{Z}$. Consequently, $f_{c}$ can be divided into two parts: the feature extraction component $z_{c}:X\rightarrow\mathcal{Z}$ and the linear classification component $l_{c}:\mathcal{Z}\rightarrow C$, where $f_{c}=z_{c}\circ l_{c}$. This means that the classification result of $f_{c}$ is achieved by first applying $z_{c}$, followed by $l_{c}$. By constraining the distribution of the mixed sample $x_{m,k}$ in the latent space to cluster around the centroid $Mean(y_{k})$ of the original category $y_{k}$, we can significantly reduce the dispersion of the noise trigger features. Thus, we can derive the latent space loss as Eq. (6), where $\mathcal{L}_{1}$ denotes the L1 loss.
@@ 4.3. FMBA Paradigm
[…]
@@
[…]
@@
[…]
The variation of FFCBA’s ASR with poisoning rate.
@@ Results
5. Evaluation
5.1. Experimental Setup
Hyperparameters. For training each classification model, we use the momentum Stochastic Gradient Descent (SGD) optimizer with an initial learning rate of 0.001 and a momentum of 0.9, decaying the learning rate by a factor of 0.1 every 30 epochs. For class-conditional autoencoders, we use the Adam optimizer with a learning rate of 0.0001. Besides, we set the PSNR threshold to 35 and the $l_{\infty}$ norm constraint to 80. The parameters $k$, $\alpha$, $\beta$, and $\gamma$ are set to 1.5, 0.5, 0.3, and 0.5, respectively. For CIFAR10, Animals90, and ImageNet100, the poisoning rate is uniformly set at 0.4%. For each baseline, we replicated its optimal experimental setup to ensure successful results.
@@ 5.2. Effectiveness of FFCBA
Attack Effectiveness. Table 2 compares FFCBA’s attack performance with baselines. FSBA and FMBA, using class-conditional autoencoders trained on the Resnet18 proxy model, target the Resnet18 victim model and execute cross-model attacks on VGG19. We outline the average ASR across all labels, Benign classification Accuracy (BA), and the Decrease Value in BA (DV) compared to the clean model. FFCBA achieves high ASR across all labels and datasets with minimal BA impact. It significantly outperforms One-to-N in ASR while matching Marksman and Universal Backdoor effectiveness under stricter clean-label constraints. Moreover, FFCBA maintains 0.4% poisoning rate, substantially lower than Marksman (10%) and One-to-N (20%), comparable to Universal Backdoor (0.62%). Figure 4 shows FFCBA’s ASR variation with poisoning rates. In addition, FFCBA’s poisoned samples exhibit superior visual quality, as illustrated in Appendix A.1. Table 3 presents FFCBA’s ASR for each CIFAR10 label with Resnet18 victim model. FFCBA demonstrates robust attack performance in all categories. Given the large number of categories in Animals90 and ImageNet100, we omit ASR for each label individually.
@@
Cross-model Attack Capability. We further evaluate the cross-model attack capabilities of FFCBA under various conditions. Table 4 presents FFCBA’s cross-model attack performance across different datasets. Specifically, we train class-conditional autoencoders on various datasets using the Resnet18 proxy model to attack other models with significant architectural and scale differences. The results indicate that FMBA maintains strong cross-model capability across all datasets while FSBA declines with increasing architectural disparity. Table 5 presents FFCBA’s cross-model attack performance across different proxy models. Specifically, we train class-conditional autoencoders on ImageNet100 using Densenet121 and ViT_B_16 proxy models, which have significant structural differences, to launch attacks on various models. The results show that FMBA maintains superior cross-model attack capabilities regardless of the proxy model used, whereas FSBA performs weaker. Despite FSBA’s cross-model limitations, Table 6 confirms its effectiveness with known victim models: When proxy model matches victim type, FSBA achieves superior attacks rapidly across architectures and datasets. Therefore, with known victim model, FSBA can be deployed for quick and efficient attacks. With unknown victim model or challenging proxy training, FMBA can be employed for cross-model attacks. Their application scenarios are complementary.
@@
5.3. Robustness against Backdoor Defenses
We assess the robustness of FFCBA against popular backdoor defenses, including Fine-Pruning (Liu et al., 2018a), Neural Cleanse (Wang et al., 2019), STRIP (Gao et al., 2019), CBD (Zhang et al., 2023), EBBA (Gao et al., 2024a), ABL (Li et al., 2021b), and IBD-PSC (Hou et al., 2024). These defenses have proven to be effective against previous backdoor attacks.
Resistance to Fine-Pruning. Fine-Pruning removes backdoors by pruning dormant neurons. Figures 5(a) and 5(b) demonstrate the resistance of FFCBA to Fine-Pruning. Across pruning rounds, ASR decline consistently remains below BA degradation, hence Fine-Pruning cannot defend against these two attack paradigms.
Resistance to Neural Cleanse. Neural Cleanse detects backdoors by constructing reverse-engineered triggers and measuring whether anomaly metric exceed a threshold of 2. Figure 5(c) shows both FSBA and FMBA maintain sub-threshold anomaly metrics, thus Neural Cleanse cannot defend against both attack paradigms.
Resistance to STRIP. STRIP detects backdoors by perturbing inputs and identifying low-entropy outputs. Figures 5(e) and 5(f) show that the entropy distributions of clean and poisoned Animals90 samples for FSBA and FMBA are similar, evading detection. The results on the other two datasets are provided in Appendix A.4.
@@ 5.3. Robustness against Backdoor Defenses
[…]
Results of FFCBA against various defenses.
Resistance to CBD. CBD mitigates backdoors via causal learning and sample weighting on contaminated dataset. Figure 5(d) confirms FFCBA’s resistance: both FSBA and FMBA maintain high ASR, demonstrating their robustness against CBD.
Resistance to EBBA. EBBA detects backdoors by identifying labels with abnormal output energy. Figures 9(c) and 10(c) show uniform label energy distributions for FSBA and FMBA on ImageNet100 (other datasets in Appendix A.5), hence EBBA fails in detecting FFCBA.
Resistance to ABL. ABL isolates backdoor samples at the early training stage and later breaks target-class correlations to train clean models from poisoned data. Figure 5(i) shows that both FSBA and FMBA maintain high ASR, indicating they can bypass ABL.
Resistance to IBD-PSC. IBD-PSC detects backdoors by amplifying batch norm parameters to enhance poisoned prediction confidence consistency. Figure 5(j) (threshold TT=0.9) shows low AUROC and F1 scores for FSBA and FMBA, thus IBD-PSC fails against FFCBA.